### v1

# Domain Knowledge Infused Generative Models for Drug Discovery Synthetic Data

Abstract
The role of Artificial Intelligence (AI) is growing in every stage of drug development. Nevertheless, a major challenge in drug discovery AI remains: Drug pharmacokinetic (PK) and Drug-Target Interaction (DTI) datasets collected in different studies often exhibit limited overlap, creating data overlap sparsity. Thus, data curation becomes difficult, negatively impacting downstream research investigations in high-throughput screening, polypharmacy, and drug combination. We propose xImagand-DKI, a novel SMILES/Protein-to-Pharmacokinetic/DTI (SP2PKDTI) diffusion model capable of generating an array of PK and DTI target properties conditioned on SMILES and protein inputs that exhibit data overlap sparsity. We infuse additional molecular and genomic domain knowledge from the Gene Ontology (GO) and molecular fingerprints to further improve our model performance. We show that xImagand-DKI-generated synthetic PK data closely resemble real data univariate and bivariate distributions, and can adequately fill in gaps among PK and DTI datasets. As such, xImagand-DKI is a promising solution for data overlap sparsity and may improve performance for downstream drug discovery research tasks. Code available at: https://github.com/GenerativeDrugDiscovery/xImagand-DKI
1 Introduction
Artificial intelligence (AI) is set to substantially reduce the \$2-3 billion dollars and 10-15 years typically required to bring a drug candidate to market (Kim et al., 2021; Wouters et al., 2020). Fewer than 10% of drug candidates successfully reach the market (Wouters et al., 2020), with the vast majority failing in clinical development due to safety and lack of activity (Paul et al., 2010). Drug discovery fails for two main reasons (Hughes et al., 2011): lack of efficacy and safety concerns. Understanding the relationship between pharmacokinetics and drug-response is essential for effective drug development (Kawabata et al., 2011; Bhalani et al., 2022).
AI is gaining momentum in drug discovery by enabling innovative preclinical approaches, including target selection and identification (Murmu & Győrffy, 2024), drug repurposing (Thafar et al., 2022; Park & Cho, 2025), drug-target interactions (DTI) (Lian et al., 2021), drug property prediction (Kim et al., 2021), de novo generation (Vignac et al., 2023; Hu et al., 2024), and synthetic data generation (Hu et al., 2025).
These advances in AI-driven drug discovery have been fueled by ongoing efforts to promote open access to data for AI training and testing (Huang et al., 2021; Brown et al., 2019; Gaulton et al., 2017). However, sequence-based molecular and biological representations, such as SMILES and amino acid sequences, alone are likely not sufficient in fully capturing the complexity of natural entities like drug molecules, proteins, and omics data. Beyond binding affinity and target specificity, modern discovery pipelines must account for a wide range of pharmacokinetic (PK) and pharmacodynamic (PD) properties, including membrane permeability (Menichetti et al., 2019), metabolic stability, bioavailability, and toxicity (e.g., LD50).
Motivated by these advances, we present xImagand-DKI, a novel multi-view SMILES/Protein-to-PK/DTI (SP2PKDTI) diffusion model. Conditioned on SMILES and protein embeddings, xImagand-DKI is capable of simultaneously generating 9 PK properties and 3 DTI values. Our key contributions are as follows:
  • Proposes an end-to-end framework that unifies PK property prediction and DTI modeling into a single foundational model, advancing solutions to data sparsity by generating high-quality synthetic drug discovery data.
  • Introduces multi-view domain knowledge integration methods that incorporate protein knowledge from the Gene Ontology(GO) (Aleksander et al., 2023) and various molecular fingerprints
  • Demonstrates how end-to-end training method combined with multi-view domain knowledge integration can effectively address the challenge of data sparsity, bridging the gap between PK and DTI datasets.
Figure 1: Visualizing data overlap sparsity between PK datasets and between DTI datasets (top), and between PK and DTI datasets (bottom). We observe 16% of PK and 4.7% of DTI molecules with overlap. See Appendix B for additional details on data sparsity.
Notably, xImagand-DKI generates dense synthetic data that addresses the challenges posed by sparse and non-overlapping PK and DTI datasets. This fragmentation, as evident in Figure 1, poses a major barrier for researchers aiming to address complex questions that require integrated data, such as those in polypharmacy and drug combination studies. Using xImagand-DKI, researchers can generate large synthetic PK and DTI assay data across thousands of ligands, enabling the exploration of poly-pharmacy and drug combination research questions, at a fraction of the cost of conducting in vitro or in vivo PK assay panels.
2 Method
xImagand-DKI is an SP2PKDTI diffusion model conditioned on learned SMILES and protein embeddings from SMILES and protein encoder models to generate target PK properties and drug-target interaction values. xImagand-DKI resembles a typical vision transformer architecture (Dosovitskiy et al., 2021); see Figure 3. 1D patches are computed from the classifier-free guidance of SMILES and protein embeddings and concatenated with PK class tokens. Diffusion step embeddings are generated using sinusoidal position encodings (Vaswani et al., 2023). Patches are then fed alongside sinusoidal step embeddings (Ho et al., 2021) to a transformer base. We mask out missing values when computing the loss for the model, only to flow gradients and learn from non-missing PK values during training. Exponential Moving Average (EMA) (Tarvainen & Valpola, 2018) is applied to the base model during training to generate the final model used for sampling. Additional model details and hyperparameters can be found in Appendix A.
3 Experiments
Figure 2: Distributions of ligand PK properties. Blue, synthetic distributions; orange, real distributions.
PKs DTIs
C2 Li. Aq FS $P$ VD HL ClH ClM $\mathbf{K_{d}}$} $\mathbf{K_{i}}$ I50
Sygd 0.62 0.53 0.34 0.50 0.66 0.81 0.85 0.59 0.58 $\varnothing$ $\varnothing$ $\varnothing$
cGan 0.19 0.16 0.17 0.18 0.25 0.24 0.28 0.32 0.29 0.32 0.08 0.13
Imgd 0.19 0.12 0.13 0.18 0.20 0.27 0.36 0.20 0.19 0.27 0.13 0.11
Base 0.12 0.08 0.07 0.13 0.11 0.12 0.15 0.13 0.18 0.26 0.07 0.09
DKI 0.13 0.07 0.07 0.12 0.09 0.08 0.15 0.15 0.15 0.24 0.06 0.07
Table 1: Average Hellinger distance across 30 generated synthetic target property datasets for ablation experiment configurations. The best HD values for each ablation test are bolded. We compare our proposed model with and without DKI to existing benchmarks of Imagand, Syngand, and cGAN.
We evaluate synthetic data generated by our model against real data over 9 PK and 3 DTI datasets. Details about each of the 9 PK and 3 DTI datasets are provided in Appendix B. Synthetic data is evaluated against real data in terms of comparing distributions, Hellinger Distance, and Machine Learning Efficiency (MLE) (Basri et al., 2023; Hu et al., 2023a). Details on our evaluation metrics are defined in Appendix C. We compare our models to baselines of Conditional GAN (cGAN) (Mirza & Osindero, 2014), Syngand (Hu et al., 2024) , and Imagand (Hu et al., 2025) .
Figure 2 shows the distributions of PK synthetic data generated by xImagand-DKI with the real data. Computing the Hellinger distance, Table 1, we see an average of 0.11, meaning that our model produces synthetic data that closely resembles the distribution of real data. Table 1 shows that data generated from our proposed architecture more closely resembles real data compared to other models. section 3 shows the results of the DTI regression tasks using real and synthetic augmented datasets. Results of these experiments suggest that a synthetic augmented dataset has equivalent utility as real data over our 3 DTI datasets. Additional tasks will be explored in future work. xImagand-DKI has similar MLE performance compared to cGAN.
Models
Real cGAN Imgd Ours
C2 mse 0.63 0.17 0.13 0.06
R2 -3.2 -0.08 0.14 -0.13
pcc 0.35 0.34 0.43 0.35
Li. mse 0.17 0.14 0.15 0.09
R2 0.04 0.19 0.14 0.01
pcc 0.50 0.47 0.41 0.49
Aq mse 0.075 0.07 0.08 0.07
R2 0.56 0.57 0.53 0.38
pcc 0.76 0.76 0.73 0.75
[…]
R2 -2.5 -0.09 0.08 -0.22
pcc 0.38 0.42 0.39 0.39
$P$ mse 3.5 0.26 0.26 0.04
R2 -13 -0.08 -0.06 -0.05
pcc 0.10 0.23 0.22 0.10
VD mse 0.54 0.21 0.20 0.04
R2 -1.8 -0.06 -0.02 -0.07
pcc 0.23 0.31 0.30 0.21
Models
Real cGAN Imgd Ours
HL mse 0.53 0.28 0.26 0.07
R2 -1.6 -0.54 -0.28 -0.09
pcc 0.16 0.13 0.03 0.17
CH mse 1.9 0.43 0.43 0.15
R2 -4.2 -0.15 -0.20 -0.13
pcc 0.11 0.14 0.10 0.10
CM mse 0.72 0.20 0.21 0.04
R2 -2.6 -0.04 -0.04 -0.06
pcc 0.13 0.25 0.25 0.17
$K_{d}$ mse 0.11 0.11 0.11 0.11
R2 0.22 0.23 0.23 0.23
pcc 0.50 0.49 0.50 0.50
$K_{i}$ mse 0.11 0.11 0.11 0.11
R2 0.21 0.21 0.22 0.22
pcc 0.46 0.46 0.47 0.47
I50 mse 0.13 0.13 0.13 0.13
R2 0.16 0.16 0.16 0.16
pcc 0.40 0.40 0.40 0.40
Table 2: Comparing drug discovery Machine Learning Efficiency (MLE) regression performances between different models and with real train data. Mean Squared Error (mse), R-Squared (R2), and Pearson Correlation Coefficient (pcc) values are averaged over 30 trials, with the best scores on the real testset bolded. R2 and pcc values are scale-adjusted relative to Real-Real with cGAN and Imagand results.
4 Discussions
Our work is a major step towards building a new class of foundational models for drug discovery trained over a diverse range of datasets. Given the problem of data sparsity, xImagand-DKI can be utilized primarily as a in silico pre-clinical tool, aimed to reduce the costs of in vitro experiments and high-throughput screening. As a research tool, scientists can utilize our models to investigate and generate properties for novel molecules to be used for downstream PBPK simulations without costly assays. Even as an initial step, xImagand-DKI has many real-world pre-clinical applications where data sparsity and data scarcity are challenges.
Although we cover a wide variety of ADMET and DTI datasets, most of these datasets are in vitro. One of the critical challenges in drug discovery is quantitative in vitro-to-in vivo extrapolation (QIVIVE). QIVIVE is an approach that extrapolates from in vitro concentration-response data to in vivo safe exposures or to identify exposure levels causing adverse effects. For future work, we will look to extend our model to include in vivo datasets and to investigate new applications of xImagand-DKI for QIVIVE.
5 Conclusions
The SMILES/Protein to PK/DTI model xImagand-DTI generates synthetic PK and DTI target property data that closely resembles real data in univariate and for downstream tasks. xImagand-DKI provides a solution for the challenge of sparse overlapping PK and DTI target property data, allowing researchers to generate data to tackle complex research questions and for high-throughput screening. Future work will expand xImagand-DKI to categorical PK and DTI properties, and scale to more datasets and larger model sizes. In future work we will look to explore additional reparameterization tricks for diffusion, such as discrete diffusion (Austin et al., 2021), to extend our methodology to be capable of learning and generating synthetic data following categorical and Log-logistic distributions common in drug discovery datasets.
Appendix A xImagand-DKI Model
Figure 3: The xImagand-DKI architecture, training, and inference methodology. Embeddings for proteins and SMILES are generated using ProtBERT and DeBERTa, respectively. Protein knowledge infusion from the human Gene Ontology (GO) is generated using PO2Vec, and SMILES knowledge infusion from fingerprints is generated using FPFormer. The model undergoes 2.5k PK training steps and 3k DTI training steps every epoch.
A.1 Diffusion Model
Given samples from a data distribution $q(x_{0})$, we are interested in learning a model distribution $p_{\theta}(x_{0})$ that approximates $q(x_{0})$ and is easy to sample from. Jonathan et al. (2020) considers the following Markov chain with Gaussian transitions parameterized by a decreasing sequence $\alpha_{1:T}\in(0,1]^{T}$:
$q(x_{1:T}|x_{0}):=\mathcal{N}(x_{1:T}|\sqrt{\alpha_{1:T}}x_{0},(1-\alpha_{1:T})\mathbf{I})$ (1)
This is called the forward process, whereas the latent variable model $p_{\theta}(x_{0:T})$ is the generative process, approximating the reverse process $q(x_{t-1}|x_{t})$. The forward process of $x_{t}$ can be expressed as a linear combination of $x_{0}$ and noise variable $\epsilon$:
$x_{t}=\sqrt{\alpha_{t}}x_{0}\ +\ \sqrt{1-\alpha_{t}}\epsilon$ (2)
We train with the simplified objective:
[…]
Where $\epsilon_{\theta}:=\{\epsilon_{\theta}^{(t)}\}_{t=1}^{T}$ is a set of T functions, indexed by t, each with trainable parameters $\theta^{(t)}$.
A.2 Pre-trained SMILES and Protein Encoders
SP2PKDTI diffusion models need powerful semantic SMILE and Protein encoders to capture the complexity of arbitrary chemical and biological structure inputs. Given the sparsity and small size of PK datasets, encoders trained on specific SMILES-Pharmacokinetic or SMILES-Protein pairs are infeasible (Huang et al., 2021). Many transformer-based foundational models such as ChemBERTa (Chithrananda et al., 2020; Ahmad et al., 2022), SMILES-BERT (Wang et al., 2019), and MOLGPT (Bagal et al., 2021) have been pre-trained to deeply understand molecular and chemical structures and properties. Similar transformer-based foundation models such as ProtBERT (Elnaggar et al., 2020) have been pre-trained to deeply understand protein structures and properties. After pre-training, these foundational models can then be fine-tuned for various downstream molecular and protein tasks. Language models trained on a SMILES-only or protein-only corpus, significantly larger than the SMILES-Pharmacokinetic and SMILES-protein data, learn a richer and wider distribution of molecular, chemical, and protein structures.
We test SMILES embeddings from ChemBERTa (Ahmad et al., 2022) and protein embeddings from ProtBERT (Elnaggar et al., 2020) trained on SMILES-only and protein-only corpora, respectively. Both embedding models were collected through the Huggingface (Wolf et al., 2020) Model Hub. Similar to Saharia et al. (2022), we freeze the weights of our embedding models. Because embeddings are computed offline, freezing the weights minimizes computation and memory footprint for embeddings during model training.
A.3 Drug Discovery Domain Knowledge
The GO is one of the most widely used resources in bioinformatics, offering structured annotations that describe the functions of genes and proteins across species. However, despite its biological richness, GO has rarely been directly integrated into deep learning models for drug discovery tasks. This underutilization stems partly from the dominance of sequence-based representations, which, although effective, often fail to capture the functional hierarchies and semantic relationships encoded in GO. Motivated by this limitation, we aim to enhance the quality of target protein embeddings by incorporating ontology-based information alongside sequence-level features.
Molecular fingerprints are bit strings that encode the structural information of a molecule, such as the presence or absence of specific chemical groups, atom types, or topological features (Hu et al., 2023b). Molecular fingerprints offer a versatile representation where different algorithms tailored to capture different aspects of molecular structure, such as key-based fingerprints and hash fingerprints. Key-based fingerprints, including MACCS (Durant et al., 2002) and RDKit (Landrum, 2013), utilize a predefined fragment library to encode each molecule into a binary bit stream according to its substructure. Hash-based fingerprints such as Morgan fingerprints (Morgan, 1965) encode substructures in a molecule based on paths around atoms in a molecule. Leveraging fingerprints alongside SMILES representations in parallel increases the generalizability of models (Schimunek et al., 2023).
Appendix B Pharmacokinetic and Drug-Target Interaction Datasets
Dataset Caco. Lipo. AqSo. Free. PPBR VDss Half Cl H Cl M
DTI Overlap 338 2789 1189 112 1241 184 486 698 794
PK Overlap 179 1751 884 527 1296 163 337 879 1018
Dataset Size 906 4200 9982 642 1797 1111 665 1020 1102
Table 3: Number of overlapping molecules for each 9 PK dataset with DTI and other PK datasets. We observe that there is a greater number of unique molecules in PK datasets that overlap with DTI datasets compared to other PK datasets.
All 9 PK and 3 DTI datasets are collected from TDCommons (Huang et al., 2021). Analyzing the overlap of 9 PK and 3 DTI datasets used in this study, Table 3 reveals minimal overlap and significant sparsity across datasets. We select PK datasets suitable for regression from the absorption, distribution, metabolism, and excretion (ADME) and Toxicity categories. We select DTI datasets from BindingDB (Liu et al., 2007) covering properties such as inhibition constant ($K_{i}$), dissociation constant ($K_{d}$), and half maximal inhibitory concentration (IC50). Revealing the overlap sparsity between DTI and PK, out of around 700k molecules from BindingDB, only around 5k molecules (0.7%) have PK properties defined from one of the 11 PK datasets.
The inhibition constant ($K_{i}$) is a measure of how strongly an inhibitor binds to an enzyme, effectively indicating the inhibitor’s potency. BindingDB has 375k pairs of $K_{i}$ values from 175k drugs and 3k proteins. The dissociation constant quantifies binding affinity between a drug and its target protein, defined as the free ligand concentration at which 50% of the protein binding sites are occupied at equilibrium. BindingDB has 52k pairs of $K_{d}$ values from 11k drugs and 1.5k proteins. The half maximal inhibitory concentration (IC50) is a measure of the potency of a substance in inhibiting a specific biological or biochemical function. BindingDB has 991k pairs of IC50 values from 550k drugs and 5k proteins.
Caco-2 (Wang et al., 2016) is an absorption dataset containing rates of 906 drugs passing through the Caco-2 cells, approximating the rate at which the drugs permeate through the human intestinal tissue. Lipophilicity (Wu et al., 2018) is an absorption dataset that measures the ability of 4,200 drugs to dissolve in a lipid (e.g. fats, oils) environment. AqSolDB (Sorkun et al., 2019) is an absorption dataset that measures the ability of 9,982 drugs to dissolve in water. FreeSolv (Mobley & Guthrie, 2014) is an absorption dataset that measures the experimental and calculated hydration-free energy of 642 drugs in water.
Plasma Protein Binding Rate (PPBR) (Wenlock & Tomkinson, 2016) is a distribution dataset of percentages for 1,614 drugs on how they bind to plasma proteins in the blood. Volume of Distribution at steady state (VDss) (Lombardo & Jing, 2016) is a distribution dataset that measures the degree of concentration for 1,130 drugs in body tissue compared to their concentration in blood.
Half Life (Obach et al., 2008) is an excretion dataset for 667 drugs on the duration for the concentration of the drug in the body to be reduced by half. Clearance (Di et al., 2012) is an excretion dataset for around 1,050 drugs on two clearance experiment types, microsome and hepatocyte. Drug clearance is defined as the volume of plasma cleared of a drug over a specified time (Huang et al., 2021).
Appendix C Evaluation Metrics
C.1 Hellinger Distance
Hellinger distance (HD) quantifies the similarity between two probability distributions and can be used as a summary statistic of differences for each PK target property between real and synthetic datasets. Given two discrete probability distributions $P=\{p_{1},p_{2},...,p_{n}\}$ and $Q=\{q_{1},q_{2},...,q_{n}\}$, the HD between $P$ and $Q$ is expressed in Equation 4.
$HD^{2}(p,q)=\frac{1}{2}\sum^{n}_{i=1}\left(\sqrt{p_{i}}-\sqrt{q_{i}}\right)$ (4)
With scores ranging between 0 to 1, HD values closer to 0 indicate smaller differences between real and synthetic data and are thus desirable.
C.2 Machine Learning Efficiency
Machine Learning Efficiency (MLE) is a measure that assesses the ability of the synthetic data to replicate a specific use case (Dankar & Ibrahim, 2021; Basri et al., 2023; Borisov et al., 2022). MLE represents the ability of the synthetic data to replace or augment real data in downstream use cases. To measure MLE, two models are trained separately using synthetic versus real data, and then their performance, measured by Mean-Squared Error (MSE), R-Squared (R2), and Pearson Correlation Coefficient (PCC), is evaluated on real data test sets and compared.
For this experiment, we train Linear Regression (LR) models using ChemBERTa and ProtBERT embeddings to predict each PK and DTI target property value. To prevent data leakage, we first divide real and synthetic data before combining them to form train and test sets.